OpenStudy (anonymous):

Inverse of W(t)=200(1.1-e^-t)? I tried but got stuck...

OpenStudy (anonymous):

W(t)=200(1.1-e^-t) =220-200e^-t -220=-200e^-t 1.1=e^-t ln1.1=lne^-t ln1.1=-t t=ln(-1.1)

OpenStudy (anonymous):

-ln(1.1-.005w) is what my answer sheet sayas?

OpenStudy (anonymous):

\[Let \space y = 200(1.1-e^{-t})\] \[y = 220 - 200e^{-t}\] \[200e^{-t} = 220-y\] \[e^{-t} = \frac{(220-y)}{200}\] \[-t = \ln\left(\frac{220-y}{200}\right)\] \[-t = \ln(220-y) - \ln(200)\] \[t = \ln(200)-\ln(220-y)\] Replace all ys with t and t with W^-1(t) \[W^{-1}(t) = \ln(220) - \ln(220-y)\]

OpenStudy (anonymous):

Sorry, last line should be: \[W^{-1}(t) = \ln(200) - \ln(220-t)\]